Anne, who is a Canadian saver, would like to invest $65,000 dollars (Canadian dollars to be precise). She is considering two options, buying a Canadian discount bond or a Russian discount bond, and Anne would like you compare returns on both options. The world risk-free rate is 0.25%. There is no risk premium on the Canadian discount bond, and the risk premium on the Russian discount bond is 5%. The current nominal Russian-Canadian exchange rate is ecan=56 Rubbles (Rubbles per Canadian dollar). Before the pay-out next

period, you expect (and of course, Anne agrees with you), Rubbles will depreciate relative to Canadian dollar, efuture=60 Rubbles (again, Rubbles per Canadian dollar). Based on your forecast, what is the expected rate of return (% yield) on each investment. For simplicity, please assume zero transaction costs and no difference in taxes (zero taxes on both options).


1
Expert's answer
2020-10-19T13:36:39-0400


find the required profitability separately

let the market risk premium be 4.5%

"\beta=1"

then for the bond of Russia:

"0.25+5+4.5\times1=9.75"

then for canadian bonds

"0.25+0+4.5*1=4.75"

let the year of circulation

then for the bond of Russia:

"FV=65 000*60(1+0.0975)^1=4 280250"

"\frac{4280250}{60}=71337.5"

71 337.5-65 000=6 337.5 - profit

for canadian bonds

"FV=65 000(1+0.0475)^1=68 087.5"

68 087.5-65 000=3 087.5

6337.5>3087.5

better invest in a Russian bond